OpenStudy (matimaticas):

Prove: sin θ - sin θ•cos^2 θ = sin^3 θ

OpenStudy (mertsj):

\[\sin \theta -\sin \theta \cos ^2\theta =\sin \theta(1-\cos ^2\theta)=\sin \theta(\sin ^2\theta)=\sin ^3\theta \]

OpenStudy (matimaticas):

@Mertsj how did you get the second part?

OpenStudy (mertsj):

factored out sin theta

OpenStudy (mertsj):

\[x-xy^2=x(1-y^2)\]